Don fenced in a rectangular garden in his yard. The length of the rectangle is 5 feet longer than the width and the perimeter is 62 feet. What is the width of the rectangle?
Answer by stanbon(75887)

Don fenced in a rectangular garden in his yard. The length of the rectangle is 5 feet longer than the width and the perimeter is 62 feet. What is the width of the rectangle?
---
Perimeter = 2(L + W)
Let width = W
Then length = W+5
---
Equation:
62 = 2(W+5 + W)
31 = 2W+5
2W = 26
W = 13 ft.
-
L = 18 ft.
=============
Cheers,
Stan H.
=============